# Efficient and Intuitive Two-Phase Validation Across Multiple Models via Principal Components

### v1

@@ Abstract
Two-phase sampling offers a cost-effective way to validate error-prone measurements in observational databases or randomized trials. Inexpensive or easy-to-obtain information is collected for the entire study in Phase I. Then, a subset of patients undergoes cost-intensive validation to collect more accurate data in Phase II. Critically, any Phase I variables can be used to strategically select the Phase II subset, often enriched for a particular model of interest. However, when balancing primary and secondary analyses in the same study, competing models and priorities can result in poorly defined objectives for the most informative Phase II sampling criterion. We propose an intuitive, easy-to-use solution that balances and prioritizes explaining the largest amount of variability across all models of interest. Using principal components to succinctly summarize the inherent variability of the error-prone covariates for all models. Then, we sample patients with the most “extreme” principal components (i.e., the smallest or largest values) for validation. Through simulations and an application to data from the National Health and Nutrition Examination Survey (NHANES), we show that extreme tail sampling on the first principal component offers simultaneous efficiency gains across multiple models of interest relative to sampling for one specific model. Our proposed sampling strategy is implemented in the open-source R package, auditDesignR.
@@ 1 Introduction
Data are rarely collected for a single limited purpose in research. Observational databases, like electronic health records (EHR), compile troves of information on extensive samples, providing new opportunities to explore various novel scientific interests with a single data source. Randomized experiments, such as clinical trials, attempt to simultaneously address the primary endpoint alongside multiple secondary objectives. Moreover, data from both observational and randomized studies can require validation in their own ways.
@@ 1.1 Necessity of Validation in Observational Studies and Clinical Trials
For observational databases, data are generally routinely collected with other purposes (e.g., providing good clinical care) at the forefront. They may have errors resulting from a variety of sources, such as improper collection, transcription, or a lack of data cleaning. Thus, when these data are repurposed for research [Nordo2019], called “secondary uses,” they may exhibit measurement error or inaccuracies [Goldstein2025] and bias downstream analyses [giganti2020]. Manual chart reviews, wherein clinical experts compare research data to clinical source documents (like patients’ electronic charts), can help to validate observational datasets. This practice has gained popularity across disciplines, including global HIV research [duda2012, Lotspeich2023] and learning health systems [shepherd2022, lotspeich2025overcomingdatachallengesenriched], but it is usually implausible to conduct chart reviews for entire studies. Partial validation (i.e., of a subset of patients) is more feasible.
@@
Source document verification (data validation) has long been standard in clinical trials [weiss1998], making data collected therein of generally high quality in terms of typos and other human errors. When data quality concerns are identified, there are methods to incorporate these audit findings into statistical models (e.g., [shepherd2012]). Further, under random assignment, the exposure of interest (generally a treatment) should be perfectly collected. Still, trial outcomes and biomarkers must be measured, and procedures and tests for them may be imperfect. Thus, validation in clinical trials can involve the re-measurement using more accurate instruments to account for measurement error [lyles2011, tang2015, Han2020]. Again, the more accurate measures are usually invasive or expensive, making them unreasonable to obtain for all trial participants.
@@
For either observational or randomized studies, validation is necessary, and partial validation is often the most feasible option, given logistical constraints like time and funding. Partial validation studies are a special type of two-phase design [white1982two]. In Phase I, inexpensive or readily available information is collected for all $N$ patients in the study (e.g., from EHR or using less-accurate, less-invasive diagnostics). In Phase II, a subset of $n$ patients ($n<N$) is chosen for validation, such that more accurate measurements are collected (e.g., from expert chart reviews or more-accurate, more-invasive testing). The sampling strategy for Phase II can leverage any information available for all patients in Phase I.
1.2 Informative Sampling Strategies for Partial Validation
@@
Often, the Phase II subset is enriched for a particular model (e.g., to minimize the variance of a model coefficient of interest). This enrichment depends on the type of model and the estimation procedure (e.g., linear regression using multiple imputation to overcome missing validation data for the rest of Phase I). Essentially, the identification of which patients would be most informative to validate depends on what we plan to analyze about them and how.
@@
For design-based estimators, like inverse probability weighting (IPW) and mean score, a variety of informative sampling designs exist for partial validation in a two-phase schema to minimize the variances of a particular model coefficient. Model-based estimators, like multiple imputation and maximum likelihood estimation (MLE), allow for more design flexibility than design-based estimators, like inverse probability weighting (IPW), since patients can be excluded from consideration (i.e., given zero validation sampling probability). Their variances do not lend themselves to intuitive, analytical solutions for optimal designs, but many have been proposed (e.g., [mcisaac2014response, tao2020optimal, amorim2021two, lotspeich2023optimal]). While not theoretically optimal, sampling the patients with extreme values of their outcome (called outcome dependent or extreme tail sampling [ETS]) is typically more efficient than simple random sampling (SRS) [feldt1961use, alf1975use, risch1995extreme, allison1997transmission, Zhou2007, lin2013quantitative]. ETS on other quantities, including (weighted) residuals [lotspeich2025overcomingdatachallengesenriched, tao2020optimal, MaoCook2023] and influence functions [amorim2021two], can offer sizable efficiency gains, as well.
@@ 1.3 Balancing Multiple Models and Objectives
Most informative two-phase sampling designs assume that estimating a single coefficient well (e.g., with great precision) is the study’s priority. Yet, post-hoc and exploratory analyses can result in competing objectives when designing a validation study. Balancing multiple analytical goals leads to an ill-defined optimization problem. Instead of a two-phase design with a clear objective, having multiple models and competing objectives complicates the optimal approach for validation sampling. Of course, a “silver bullet” would maximize all objectives (unrealistic in practice), whereas maximizing just one leads to a sub-optimal system solution. For instance, focusing on a single objective may give us good performance for that model at the detriment of all others. Various approaches exist to the multi-objective problem, but none have been applied for two-phase designs [deb2016multi].
We forgo the existing complex approaches, which are dependent upon functions of the moments of the coefficients of interest (often unknowable in practice) across multiple models, for a simpler approach. Principal component analysis (PCA), a popular nonparametric dimension reduction technique, offers a promising way to adapt existing informative sampling designs to accommodate multiple modeling objectives. The components succinctly summarize the variability within the data, and, conveniently, applying PCA on centered and scaled data (or on the data’s covariance matrix) inherently resolves the issue of differences in the variables’ scales. We explore how PCA can be used to capture information across multiple error-prone variables and guide targeted sampling strategies for partial validation.
@@ 1.4 National Health and Nutrition Examination Survey
To motivate the use of our proposed PCA-driven sampling strategy in practice, we consider the accuracy of food recall logs and dietary measurements in the National Health and Nutrition Examination Survey (NHANES) [nhanes_cdc]. For the 2021–2023 cohort, two dietary recalls were conducted via phone interviews; individuals reported the foods and drinks they had consumed within the past $24$ hours.
While convenient, even the gold standard method of food recall logs is subject to recall bias, under- or over-reporting, and fallibility of portion size estimation (see [THOMPSON20175] and references therein). The additional step of converting food intake into nutrients can be complicated by product differences in nutrient density when individuals oversimplify food items (e.g., the amount of protein in “yogurt”). We estimate the relationship between nutrient intake and relevant health outcomes, incorporating partial validation data.
@@ 1.5 Overview
In this paper, we propose a new sampling strategy for partial validation studies (two-phase designs) that balances multiple modeling objectives simultaneously. Validation sampling depends on the first principal component of all observed covariates, and the outcome variables for the multiple models may be shared or unique. Our design is simple to implement and balances competing analytical objectives in an intuitive and easy-to-understand approach. We compare the efficiency (i.e., the inverse of the variance) of our design under various measurement error scenarios and find it outperforms both SRS and ETS on a single covariate. In Section 2, we outline our proposed validation study design and chosen estimator. In Section 3, we present a robust collection of simulations showing the performance of our sampling strategy. In Section 4, we consider our data application in the NHANES study. In Section 5, we conclude with a discussion on the practicality and limitations of our approach.
@@ 2.1 Model and Data
Suppose we are conducting a study of $N$ patients on whom $P$ separate statistical models are of interest. For each model $j$, we have a continuous outcome $Y_{ji}$, an exposure $X_{ji}$, and a vector of other confounders ${\mathbf{Z}}_{ji}$ ($j\in\{1,\dots,P\}$, $i\in\{1,\ldots,N\}$). They are assumed to be related through the usual normal linear regression model:
@@
where the random errors $\epsilon_{ji}$ are independent and identically distributed following a normal distribution with mean =0=0 and variance $\sigma_{j}^{2}$. We are most interested in estimating $\beta_{1j}$ from each model, but all the exposures interest $\boldsymbol{X}_{i}=(X_{i1},\dots,X_{iP})^{\top}$ are difficult to measure.
Instead, we observe error-prone versions of these exposures $\boldsymbol{X}^{*}_{i}=\boldsymbol{X}_{i}+\boldsymbol{U}_{i}$ for all $N$ patients, where $\boldsymbol{U}_{i}=(U_{i1},\dots,U_{iP})^{\top}$ is the vector of additive measurement errors for all $P$ models. We are planning a partial validation study, such that $n$ patients ($n<N$) will have their true covariates $\boldsymbol{X}_{i}$ measured. Let $R_{i}=1$ if the $i$th subject is validated (i.e., $\boldsymbol{X}_{i}$ is observed), and $R_{i}=0$ otherwise. Under the two-phase design framework, we control the assignment of the validation indicator $R_{i}$. To satisfy the missing at random (MAR) assumption [Little1992] for our ultimate analyses, $R_{i}$ can depend only on information that is fully observed (i.e., the outcomes, error-prone exposures, and confounders). Without loss of generality, we will assume that if $R_{i}=1$, then $\boldsymbol{X}_{i}$ is fully observed (i.e., the covariates are fully validated).
We want to use these data to estimate the $P$ separate models, all of which have equal priority in this study. In Section 2.2, we discuss various validation sampling designs to select the subset of $n$ patients for $\boldsymbol{X}_{i}$ collection. In Section 2.3, we outline how multiple imputation can be used to estimate the models after only partially validating the study data (i.e., overcoming missing $\boldsymbol{X}_{i}$ for the NnN-n unvalidated patients). Throughout, our priority is to strategically design the validation study to promote the efficiency of the coefficient estimates $\hat{{\boldsymbol{\beta}}}_{1}=(\hat{\beta}_{11},\dots,\hat{\beta}_{1P})^{\top}$, obtained via multiple imputation, accompanying the exposures $\boldsymbol{X}_{i}$.
@@ 2.2 Validation Sampling Designs
Traditionally, targeted validation sampling approaches focus on a single model $j$, and the goal is to identify which $n$ patients, if validated, would minimize the variance of $\hat{\beta}_{1j}$ corresponding to the covariate $X_{ji}$ in that model (e.g., [shepherd2022, lotspeich2025overcomingdatachallengesenriched, Han2020, mcisaac2014response, tao2020optimal, amorim2021two, lotspeich2023optimal, zhou2002semiparametric]). While effective for improving the estimation of $\hat{\beta}_{1j}$, as designed, these strategies do not consider the estimation of the coefficients from the other models of interest.
Many cited works suggest that optimal designs for multiple coefficients are possible with appropriate optimality criteria, like D-optimality (e.g., minimizing the determinant of the variance-covariance matrix). However, most of these designs are already complex for a single coefficient; extending them to consider multiple simultaneously would be challenging. Therefore, we propose a novel strategy for partial validation in a multi-objective study based on the ETS design (already a common strategy for single-objective studies) and PCA.
@@
In exploratory analyses or feature creation, PCA is a common approach to dimension reduction. It seeks to explain variability in the original data through a smaller subset of newly created variables (called principal components or loadings). This property of explaining variability in a reduced dimension motivates our use of PCA in designing a validation study. We apply PCA to all the models’ error-prone exposures $\boldsymbol{X}^{*}$ to create a single variable that captures the variability between them. In particular, we take the first principal component $PC_{1}^{*}$, which will always explain the most variability in $\boldsymbol{X}^{*}$ out of all the estimated principal components. Then, we prioritize patients for validation based on $PC_{1}^{*}$, rather than focusing only on a single $X_{j}^{*}$ or trying to balance all of $\boldsymbol{X}^{*}$ simultaneously. This strategy balances the reduction in variance across all $P$ models without adding computational burden, and it seamlessly integrates with existing single-objective validation sampling methods, like ETS.
@@
In our simulations and application to NHANES, we consider three sampling strategies. Simple random sampling (SRS): The $n$ patients are chosen from the original study of $N$ completely at random (i.e., with equal probability). This design ignores all available information, making $\boldsymbol{X}_{i}$ missing completely at random [Little1992]. Extreme tail sampling on the $j$th model’s error-prone exposure (ETS-$X_{j}^{*}$): The original $N$ patients are ordered ascendingly by $X_{j}^{*}$, and those with the $n/2$ largest and $n/2$ smallest values are chosen for validation. In practice, the model $j$ would be chosen a priori. For demonstration, we use $j=1$. Extreme tail sampling on the first principal component of all error-prone exposures (ETS-$PC_{1}^{*}$): The original $N$ patients are ordered ascendingly by $PC_{1}^{*}$, and those with the $n/2$ largest and $n/2$ smallest values are chosen for validation.
The intuition behind ETS designs comes from the belief that extreme observations are most informative. For example, by validating patients following the ETS-$X_{1}^{*}$ design, we correct the measurement error in the most extreme error-prone exposures and expect better estimation for the linear regression of $Y_{1}$ on $X_{1}$ (controlling for ${\mathbf{Z}}_{1}$) as these extreme values may be more influential. Our ETS-$PC_{1}^{*}$ design is similarly motivated. After expressing the entirety of $\boldsymbol{X}^{*}$ in a lower dimension (as $PC_{1}^{*}$), we seek to validate patients with the most extreme values of $\boldsymbol{X}^{*}$ and improve estimation across all $P$ linear regression models.
The advantages of the ETS-$PC_{1}^{*}$ design hinge upon how much variability in $\boldsymbol{X}^{*}$ is captured by $PC_{1}^{*}$. When the error-prone exposures $\boldsymbol{X}^{*}$ are correlated, $PC_{1}^{*}$ should perform well. Accordingly, we anticipate that the average variance across $\hat{\beta}_{1}$ should be lower using ETS-$PC_{1}^{*}$ than ETS-$X_{j}^{*}$ or SRS. However, even with uncorrelated $\boldsymbol{X}^{*}$, most models’ estimates will have similar variability with $PC_{1}^{*}, ETS$-$X_{j}^{*}$, or SRS. (The one exception would be the $j$th model, which would see much better gains under an ETS-$X_{j}^{*}$ design.)

### v2

# Two-phase validation sampling via principal components to improve efficiency in multi-model estimation from error-prone biomedical databases

@@ Abstract
Two-phase sampling offers a cost-effective way to validate error-prone covariate measurements in biomedical databases. Inexpensive or easy-to-obtain information is collected for the entire study in Phase I. Then, a subset of patients undergoes cost-intensive validation (e.g., expert chart review) to collect more accurate data in Phase II. Critically, any Phase I variables can be used to strategically select Phase II, often enriched for a particular model of interest, and then missing data methods like multiple imputation can be used for estimation. However, when balancing primary and secondary analyses, competing models and priorities can result in poorly defined objectives for the most informative Phase II sampling criterion. Extreme tail sampling (ETS), wherein patients with the smallest and largest values of a particular quantity (like a covariate or residual) are selected, can offer great statistical efficiency in two-phase studies when focusing on a single analytic objective by targeting observations with the biggest contributions to the Fisher information. We propose an intuitive, easy-to-use approach that extends ETS to balance and prioritize explaining the largest amount of variability across multiple models of interest. Using principal components, we succinctly summarize the inherent variability of all models’ error-prone exposures. Then, we sample patients with the most extreme principal components for validation. Through simulations and an application to the National Health and Nutrition Examination Survey (NHANES), the proposed sampling strategy offered simultaneous efficiency gains across multiple models of interest. Its advantages persisted across various real-world scenarios: different covariance structures, measurement error severities, validation proportions, and shared versus distinct model outcomes. When designing a validation study, concentrating on a single model may be short-sighted. Allocating resources more broadly, while still strategically, could balance multiple analytical goals simultaneously. Employing dimension reduction before sampling will allow this strategy to scale up well to big-data applications with many error-prone covariates.
@@ 1 Introduction
Data are rarely collected for a single purpose in research. Observational databases, like electronic health records (EHR), compile troves of information on extensive samples, providing new opportunities to explore various novel scientific interests with a single data source. Randomized experiments, such as clinical trials, attempt to simultaneously address the primary endpoint alongside multiple secondary objectives. Moreover, data from both observational and randomized studies require validation in their own ways.
@@ 1.1 Necessity of Validation in Observational Studies and Clinical Trials
Observational data are generally routinely collected with other purposes (e.g., providing good clinical care) at the forefront. They may have errors resulting from various sources, like improper collection, transcription, or data cleaning. Thus, when repurposed for research [28], these data may exhibit measurement error [14] and bias downstream analyses [13]. Manual chart reviews, wherein clinical experts compare research data to clinical source documents (like patients’ electronic charts), can validate observational datasets (e.g., [9, 19, 21, 34]), but they are usually implausible for entire studies. Partial validation (i.e., of a subset of patients) is more feasible.
@@
Source document verification (data validation) has long been standard in clinical trials [42], making data collected therein of generally high quality in terms of typos and other human errors. Further, under random assignment, the exposure of interest (generally a treatment) should be perfectly collected. Still, other variables, like biomarkers, may be imperfectly measured. Validation in the form of re-measurement using more accurate instruments can account for errors [15, 22, 35], but the more accurate measures are usually invasive or expensive, making them unreasonable to obtain for all trial participants.
1.2 Informative Validation Sampling Strategies
@@
Partial validation studies are a special type of two-phase design [43]. In Phase I, inexpensive or readily available information is collected for all patients in the study. In Phase II, a subset of patients is chosen for validation, such that more accurate measurements are collected. Sampling for Phase II can leverage any information available for all patients in Phase I, opening the door to many targeted strategies.
@@
Often, the Phase II subset is enriched for a particular model (e.g., to minimize the variance of a model coefficient of interest). This enrichment depends on the model type and estimation procedure. Essentially, identifing which patients would be most informative to validate depends on what we plan to analyze about them and how.
@@
For design-based estimators, like inverse probability weighting (IPW) and mean score, multiple informative sampling designs exist to minimize the variances of a particular model coefficient. Model-based estimators, like multiple imputation and maximum likelihood estimation (MLE), allow for more design flexibility than design-based estimators, since some patients can be excluded from consideration (i.e., given zero validation sampling probability). While the variances of model-based estimators do not lend themselves to intuitive, analytical solutions for optimal designs (as with design-based ones), many have been proposed (e.g., [3, 20, 24, 37]).
When no covariate information is available pre-validation, sampling patients with extreme outcome values (called outcome dependent or extreme tail sampling [ETS]) is typically more efficient than simple random sampling (SRS) [1, 2, 11, 17, 29, 31, 45]. When covariate information is available, other ETS strategies, including sampling based on (weighted) residuals [8, 21, 23, 37] and influence functions [3], can offer sizable efficiency gains, as well.
@@ 1.3 Balancing Multiple Models and Objectives
Most informative two-phase sampling designs prioritize estimating a single coefficient well (e.g., with great precision). Yet, post-hoc and exploratory analyses can result in competing objectives when designing a validation study. Balancing multiple analytical goals across various models leads to an ill-defined optimization problem, which complicates the optimal approach for validation sampling. Of course, a “silver bullet” would maximize all objectives (unrealistic in practice), whereas maximizing just one leads to a sub-optimal system solution. For instance, focusing on a single objective may give us good performance for that model at the detriment of all others. Various approaches exist to the multi-objective problem [7], but, to our knowledge, none have been applied for two-phase designs.
Optimally designing a validation study to efficiently estimate multiple coefficients in the same model poses a related, yet unique challenge. This setting, wherein a single model is of interest and high statistical precision for many or all coefficients is desired, has received more attention in the two-phase literature than ours with multiple models. For example, Sauer et al. (2021) [31] demonstrate the trade-off between focusing a two-phase design strategy on one versus many model coefficients. Considering optimal allocation strategies that minimize (i) the variance of a single coefficient and (ii) the trace of the coefficients’ variance-covariance matrix (i.e., the sum of their variances), they show that while the latter offers smaller efficiency gains for any particular coefficient it offers better precision for the model as a whole.
Intuitively, we expect the same trade-off when estimating coefficients across multiple models: We can be great for one coefficient and poor for the rest, or we can be good for all of them. While extending existing optimal designs from one model to many is possible, many depend upon functions of the moments of the coefficients of interest (often unknowable in practice). Herein, we propose a simpler approach that scales up well with the number of models and variables, making it particularly appealing for big-data applications.
Principal component analysis (PCA), a popular nonparametric dimension reduction technique, offers a promising way to adapt existing informative sampling designs to accommodate multiple modeling objectives. The principal components succinctly summarize the variability within the data, and, conveniently, applying PCA on centered and scaled data or on the data’s covariance matrix inherently resolves the issue of differences in the variables’ scales. We explore how PCA can capture information across multiple error-prone variables, each of interest in its own statistical model, and guide targeted sampling strategies for partial validation.
@@ 1.4 National Health and Nutrition Examination Survey
We consider the accuracy of food recall logs and dietary measurements in the National Health and Nutrition Examination Survey (NHANES) [12]. For the 2021–2023 cohort, two dietary recalls were conducted via phone interviews. Individuals reported the foods and drinks they had consumed within the past $24$ hours.
While convenient, even the gold standard method of food recall logs is subject to recall bias, under- or over-reporting, and fallibility of portion size estimation (see [38] and references therein). Converting food intake into nutrients can be further complicated by product differences in nutrient density when individuals oversimplify food items (e.g., the amount of protein in “yogurt”). We estimate the relationship between error-prone nutrient intake and relevant health outcomes, comparing statistical precision under various sampling strategies.
@@ 1.5 Overview
We propose a new informative sampling strategy that balances multiple modeling objectives simultaneously. Validation sampling depends on the first principal component of all observed error-prone exposures, and the outcome variables for the multiple models may be shared or unique. Our design is simple to implement and balances competing analytical objectives in an intuitive and easy-to-understand approach. The resulting partially validated data are analyzed using multiple imputation, but other model-based methods could be used. In simulations and an application to error-prone, self-reported food intake data from the NHANES study, we show that the proposed design outperforms both SRS and ETS focused on a single model in terms of statistical efficiency (i.e., the inverse of the estimated coefficients’ variances).
Statement of Significance
  • Problem Two-phase sampling offers a cost-effective way to partially validate error-prone measurements in biomedical databases (e.g., through expert chart review or evaluation with more precise instruments).
  • What Is Already Known Any fully-observed variables can be used to strategically select the validation subset, which is often enriched for a particular model of interest, and then missing data methods like multiple imputation can be used for estimation from partially validated data. However, nearly all such targeted validation sampling strategies promote efficiency in one or more coefficients in a single model.
  • What This Paper Adds We propose an intuitive, easy-to-use approach that balances and prioritizes explaining the largest amount of variability across multiple models of interest. First, principal components analysis is used to summarize variability in the error-prone data, and then observations with the most extreme principal components (i.e., the largest or smallest) are chosen for validation.
@@ 2.1 Model and Data
Suppose we are conducting a study of $N$ patients on whom $J$ separate statistical models are of interest. For each model $j$, we have a continuous outcome $Y_{ji}$, an exposure of interest $X_{ji}$, and a vector of other confounders ${\mathbf{Z}}_{ji}$ ($j\in\{1,\dots,J\}$, $i\in\{1,\ldots,N\}$). These variables are assumed to be related through the usual normal linear regression model:
@@
where the random errors $\epsilon_{ji}$ are independent and identically distributed (i.i.d.) following a normal distribution with mean =0=0 and variance $\sigma_{j}^{2}$. We are most interested in estimating $\beta_{1j}$ from each model, but all the exposures $\boldsymbol{X}_{i}=(X_{1i},\dots,X_{Ji})^{\top}$ are difficult to measure accurately.
Instead, we observe error-prone versions $\boldsymbol{X}^{*}_{i}=\boldsymbol{X}_{i}+\boldsymbol{U}_{i}$ for all $N$ patients, where $\boldsymbol{U}_{i}=(U_{1i},\dots,U_{Ji})^{\top}$ is the vector of additive measurement errors for patient $i$’s exposures from all $J$ models. We are planning a partial validation study, such that a subset of $n$ patients ($n<N$) will have $\boldsymbol{X}_{i}$ measured. To satisfy the missing at random (MAR) assumption [18] for our ultimate analyses, validation sampling can depend only on information that is fully observed (i.e., the outcomes, error-prone exposures, and confounders). Without loss of generality, we will assume that if patient $i$ is sampled for validation, then $\boldsymbol{X}_{i}$ is fully observed. We want to use these partially validated data to estimate the $J$ separate models, all of which have equal priority in this study.
Throughout, our priority is to strategically design the validation study to promote the efficiency of the coefficient estimates $\hat{{\boldsymbol{\beta}}}_{1}=(\hat{\beta}_{11},\dots,\hat{\beta}_{1J})^{\top}$, accompanying the exposures $\boldsymbol{X}_{i}$, obtained via multiple imputation. Other model-based estimation approaches, like MLE, should benefit from similar efficiency gains under the proposed design. We focused on multiple imputation here for its simpler implementation.
@@ 2.2 Validation Sampling Designs
Traditionally, targeted sampling focuses on one model $j$, and the goal is to identify which $n$ patients, if validated, would minimize the variance of $\hat{\beta}_{1j}$ corresponding to the exposure of interest $X_{ji}$ [3, 15, 20, 21, 24, 34, 37, 46]. Designs have also been proposed to improve estimation across multiple coefficients ($\hat{\beta}_{1j},\hat{\boldsymbol{\beta}}_{2j}$) in a given model by minimizing the trace (A-optimality) or determinant (D-optimality) of their variance-covariance matrix [10, 31]. However, these strategies do not consider the estimation of the coefficients from the other J1J-1 models.
Most of these designs are already complex for a single coefficient from one chosen model; extending them to consider multiple coefficients simultaneously, potentially each with its own outcome, would be challenging. Further, theoretically optimal designs to estimate $\hat{\beta}_{1j}$ are functions of true $\beta_{1j}$. Multi-wave strategies to approximate theoretically optimal designs are becoming commonplace [20, 21, 25, 34], but they can be more logistically challenging. Therefore, we instead propose a novel strategy for partial validation in a multi-objective study based on the ETS design (already a common strategy for single-objective studies) and using PCA to create a single measure summarizing variability across all models’ error-prone exposures.
@@
In exploratory analyses or feature creation, PCA is a common approach to dimension reduction. Through a smaller subset of newly created variables (called principal components or loadings), PCA seeks to explain variability in the original data. This property is central to our use of PCA in validation study design. We apply PCA to all the models’ error-prone exposures $\boldsymbol{X}^{*}$ (and, if desired, additional confounders) to create a single variable capturing the variability between them. In particular, we take the first principal component $PC_{1}^{*}$, which will always explain the most variability out of all the estimated principal components. Then, we prioritize patients for validation based on $PC_{1}^{*}$, rather than focusing only on a single $X_{j}^{*}$ or trying to balance all of $\boldsymbol{X}^{*}$ simultaneously. This strategy balances the reduction in variance across all $J$ models without adding computational burden, and it seamlessly integrates with existing single-objective validation sampling methods, like ETS.
@@
Throughout this paper, we consider three sampling strategies for the validation of $n$ patients out of a study of size $N$. As the baseline design, we include simple random sampling (SRS). The $n$ patients are chosen with equal probability ($n/N$), which ignores all available information and makes $\boldsymbol{X}$ missing completely at random (MCAR) [18]. To represent a typical existing strategy, wherein the validation study is enriched for the model of primary interest, we include extreme tail sampling on the primary model’s error-prone exposure (ETS-$X_{p}^{*}$). The original $N$ patients are ordered ascendingly by $X_{p}^{*}$ ($p\in\{1,\dots,J\}$), and those with the $n/2$ largest and $n/2$ smallest values are chosen for validation. In practice, the primary model would be chosen a priori, for example, based on clinical interest. Finally, we implement our proposed strategy, extreme tail sampling on the first principal component of all error-prone exposures (ETS-$PC_{1}^{*}$). The original $N$ patients are ordered ascendingly by $PC_{1}^{*}$, the first principal component derived from $\boldsymbol{X}^{*}$ for all $J$ models, and those with the $n/2$ largest and $n/2$ smallest values are chosen for validation.
2.3 Intuition Behind Extreme Tail Sampling
The intuition behind ETS designs comes from the belief that extreme observations (i.e., those in the tails of the distribution) are most informative to conducting statistically efficient analyses. The statistical underpinnings behind this intuition stem from the large-sample theory behind MLE (the method by which common statistical models are fit).
2.3.1 For the Primary Model
Recall that the primary model of interest captures the relationship between $Y_{p}$ and $X_{p}$ given ${\mathbf{Z}}_{p}$. In linear regression, observations with covariate values $X_{p}$ that are farthest from the mean $\overline{X}_{p}$ make the largest contributions to the Fisher information for the corresponding model coefficient $\beta_{1p}$ and help to reduce variability in its estimate $\hat{\beta}_{1p}$. Thus, these extreme values of $X_{p}$ are deemed “most informative” when estimating our model, and, if can only measure the exposure for some subset of $n$ patients, focusing on those the smallest and largest values should offer the best statistical precision. (This fourth implausible strategy could be denoted by ETS-$X_{p}$.)
However, we cannot prioritize patients for validation based on their values of $X_{p}$ (which is attainable only through validation), and the ETS-$X_{p}^{*}$ design is perhaps the most natural alternative. The hope is that by sampling on the extremes of the error-prone version $X_{p}^{*}$, we can capture many of the same patients as we would have under the ETS-$X_{p}$ design. The efficiency advantages of the ETS-$X_{p}^{*}$ relative to the ETS-$X_{p}$ will depend on how informative $X_{p}^{*}$ is about $X_{p}$. See Figure 1 for an illustration, wherein we compare which patients would be selected for validation based on sampling from the extremes of the true versus error-prone exposure under increasing error severities.
Figure 1: Plot of the primary model’s true exposure $X_{p}$ against the error-prone version $X_{p}^{*}=X_{p}+U$ under varied error severities, driven by the variance of the errors $U$. Each plot contains $N=1000$ simulated observations, and the points are colored by their validation status based on extreme tail sampling on $X_{p}^{*}$ (ETS-$X_{p}^{*}$) versus extreme tail sampling on $X_{p}$ (ETS-$X_{p}$). “Counts” denotes the number of observations in each setting with the corresponding validation status. As the error variance increases, and $X_{p}^{*}$ becomes less informative about $X_{p}$, the number of observations that would be sampled by both ETS designs decreases, and the efficiency of ETS-$X_{p}^{*}$ is expected to drop below that of ETS-$X_{p}$. Still, even under the worst error setting, the ETS-$X_{p}^{*}$ design should capture more information than simple random sampling (SRS).
2.3.2 For Multiple Models
With competing analytic objectives, we shift our design focus to reducing the sum of the variances for coefficients $\hat{\beta}_{1j}$ $(j\in\{1,\ldots,J\})$ across all models. Hereafter, this sum is referred to as the total coefficient variability across all models. Improving the estimation of $\beta_{1p}$ in the primary model, through the ETS-$X_{p}^{*}$ design or otherwise, does not guarantee any improvement in the estimation of coefficients for any other models, and thus it will not help to minimize this total. As before, the most informative observations to each individual model would be those with the most extreme exposures $\boldsymbol{X}$. Still, these values are not available when designing the validation stage, so the next-best thing is the error-prone versions $\boldsymbol{X}^{*}$. However, there is now an added challenge of tailoring the design strategy to prioritize the most extreme values across multiple variables at once.
To this end, we consider how to describe the variability across across all $\boldsymbol{X}^{*}$ in a single dimension and then incorporate that information into validation sampling. Conducting principal component analysis and taking the first principal component, $PC_{1}^{*}$, we obtain a single transformed variable that describes the greatest variability in the set of exposures $\boldsymbol{X}^{*}$, providing a compact criterion on which to sample while balancing all modeling objectives. Our ETS-$PC_{1}^{*}$ design is similarly motivated as the ETS-$X_{p}^{*}$ design. By expressing the entirety of $\boldsymbol{X}^{*}$ in a lower dimension as $PC_{1}^{*}$, we seek to validate patients with the most extreme values of $\boldsymbol{X}^{*}$ (and, ideally, $\boldsymbol{X}$), thereby improving estimation efficiency across all $J$ linear regression models. In Figure 2, we broaden the simulated example from above to evaluate five models of interest.
Figure 2: Plot of all models’ true exposure $X_{j}$ against the first principal component $PC_{1}^{*}$ summarizing all the error-prone exposures $\boldsymbol{X}^{*}$, focusing on the moderate error setting (error variance =0.5=0.5). Each plot contains the same $N=1000$ simulated observations, and the points are colored by their validation status based on extreme tail sampling on the first principal component $PC_{1}^{*}$ (ETS-$PC_{1}^{*}$) versus extreme tail sampling on the specific model’s exposure $X_{j}$ (ETS-$X_{j}$). “Counts” denotes the number of observations in each setting with the corresponding validation status. Across all five models, the number of overlapping observations that would be sampled by both ETS designs is relatively stable, and the efficiency of ETS-$PC_{1}^{*}$ is expected to offer similarly good approximations to the ETS-$X_{j}$ for each model separately.
The ETS-$PC_{1}^{*}$ design’s ability to reduce the total coefficient variability across all models depends on two characteristics of the error-prone exposure data. First, as in the primary model case, the efficiency gains are driven by how much information we can glean from the validated exposures $\boldsymbol{X}$. When the errors in $\boldsymbol{X}^{*}$ are less noisy, sampling on its extremes is expected to help us capture more informative observations in the extremes of $\boldsymbol{X}$. Second, our ability to sample the most extreme observations with respect to all of $\boldsymbol{X}^{*}$ (and subsequently $\boldsymbol{X}$) depends on how much variability is explained by the first principal component $PC_{1}^{*}$. When the error-prone exposures $\boldsymbol{X}^{*}$ are highly correlated, $PC_{1}^{*}$ will be more informative and sampling on its extremes will lead to higher efficiency gains. In the conservative case where the $\boldsymbol{X}^{*}$ are completed uncorrelated, the PCA will yield a separate factor for each exposure, and the ETS-$PC_{1}^{*}$ design will be equivalent to an ETS-$X_{j}^{*}$. In practice, when the exposures are completely independent, nominal correlation between them due to random chance can result in a slight inefficiency when using the proposed design compared to sampling on a single variable.